OpenStudy (anonymous):

A rectangular garden has a perimeter of 92 feet. If the length of the garden is 1 foot greater than twice its width, what is the length of the garden in feet?

OpenStudy (anonymous):

Step 1) Make two equations. 2l + 2w = 92 l = 1 + 2w

OpenStudy (anonymous):

The perimeter of a rectangular garden is 326 feet. If the length of the garden is 95 feet, what is its width?

OpenStudy (anonymous):

Now use substitution. Substitute the value of l into the first equation. Solve for w. 2l + 2w = 92 l = 1 + 2w __ 2(1 + 2w) = 92 Distribute the 2. 2 + 4w = 92 Subtract 2. 4w = 90 Divide by 4. w = 22.5 -- Substitute the value of the width (w = 22.5) into any equation. Solve for l(length). 2l + 2w = 92 2l + 2(22.5) = 92 2l + 45 = 92 Subtract 45. 2l = 47 Divide by 2. l = 23.5 Length is 23.5

OpenStudy (anonymous):

@leebulls

OpenStudy (anonymous):

The perimeter of a rectangular garden is 326 feet. If the length of the garden is 95 feet, what is its width? P = 2l + 2w 326 = 2(95) + 2w 326 = 190 + 2w Subtract 190. 136 = 2w Divide by 2. 68 = w Width is 68.